# How to Expose a Black Hole

### 04 Apr 2026 v1

@@ 1 Introduction
The main idea will be to use the correspondence principle formulated in [1, 2, 3], following earlier suggestions in [4, 5, 6, 7]. More recent perspective on this can be found in [8]. This principle states that as we adiabatically change the string coupling from finite value to a sufficiently small value, a Schwarzschild black hole transforms to a highly excited elementary string state. This was generalized to a wide class of other charge carrying black holes, although at weak coupling these black holes correspond to systems of D-branes and fundamental strings instead of just fundamental strings.
@@ 2.3 Varying dilaton background from charged black hole
Using these results we can also test the other conditions. For example using the lower bound on $C$ given in (2.32) we can verify that the mass of the background black hole, which is of order $C$, is much larger than the mass of the rolling black hole which is of order $N^{(D-3)/\{2(D-2)\}}$. Hence the backreaction of the rolling black hole on the background black hole geometry remains small. In particular, the causal structure of space-time remains unaffected and as long as the rolling black hole stays outside the horizon of the background black hole, given by the condition $\rho^{D-3}>2m$, it will be causally connected to the asymptotic observer. Also, it follows from (2.3) that for $\bar{\rho}\sim\bar{\rho}_{0}$, any scalar constructed from the fields and two derivatives scales as $\bar{\rho}_{0}^{-2/(D-2)}$. The only exception is $F_{\mu\nu}F^{\mu\nu}$, but the combination $e^{-4\phi/(D-2)}F_{\mu\nu}F^{\mu\nu}$ that appears in the Lagrangian density still scales as $\bar{\rho}_{0}^{-2/(D-2)}$. Hence for large $\bar{\rho}_{0}$, locally the rolling black hole experiences almost flat space-time. It also follows from this that a stationary observer in this background feels acceleration of order $\bar{\rho}_{0}^{-1/(D-2)}$. In the sense described in footnote 1, this also gives an order of magnitude estimate of the temperature experienced by the rolling black hole / string. From (2.31) we see that this is much smaller than the temperature $T_{bh}\sim N^{-1/\{2(D-2)\}}$ given in (2.14). Since (2.31) ensures that the effect of evaporation remains small during the transit, it now follows that the effect of absorption of the ambient radiation also remains small during the transit.
@@ 2.4 Fall from the asymptotic region
For $D\geq 5$ and large $N$, the upper bound is larger than the lower bound, and hence we can find $\bar{\rho}_{0}$ satisfying these conditions. For $D=4$, the $N$ dependence of the two sides are identical, but by taking $\eta$ to be small, we can still ensure that the upper bound is larger than the lower bound. Thus we see that it is possible for the rolling black hole to roll all the way from a region where the string coupling is small but $N$ independent to the region where it becomes an elementary string excitation. On the other hand, if the rolling black hole undergoes a controlled fall as discussed in footnote 1, maintaining its speed at order unity but not parametrically close to unity, then the lower bound on $\bar{\rho}_{0}$ will be given by the earlier bound $N^{1/2}$, and even for $D=4$ and finite $\eta$, the upper bound on $\bar{\rho}_{0}$ remains larger than the lower bound for large $N$.
Finally we can consider the fall from the asymptotic region $\rho\sim C^{1/(D-3)}$. In this case the transit time is of order $C^{1/(D-3)}$. Comparing this with the transit time $\bar{\rho}_{1}^{1/(D-2)}\sim C^{1/(D-3)}\eta^{2/\{(D-2)(D-3)\}}$ computed from (2.36) and (2.34), we see that this corresponds to setting $\eta\sim 1$ in the previous analysis. Also from (2.3) and (2.27) we see that the redshift factor $\gamma$ at a position $\rho\sim\bar{\rho}_{0}/C$ is now of order $(C\rho^{3-D})^{(D-3)/(D-2)}\sim N^{(D-3)/\{2(D-2)\}}$. Comparing this with (2.38) we again see that that this corresponds to setting $\eta\sim 1$ in the previous analysis. Thus for $D\geq 5$, even a free fall of the rolling black hole from the asymptotic region will convert a black hole to an elementary string state in a controlled fashion. For $D=4$ the adiabatic approximation will break down near the end of the roll, but this can be avoided by arresting the fall by external force as described in footnote 1 so that the rolling black hole does not become ultra-relativistic during the roll. In the extreme case, if we bring the rolling black hole at rest at $\bar{\rho}\sim\bar{\rho}_{0}\epsilon^{2/(D-3)}$ where it becomes an elementary string state, the proper acceleration felt by the rolling black hole will be of order $\bar{\rho}_{0}^{-1/(D-2)}\epsilon^{-2/\{(D-2)(D-3)\}}$. This remains small compared to the string scale as long as (2.30) holds. An observer in the rest frame of this black hole can study its quantum state and send the information back to the asymptotic observer since both the rolling black hole and the observer will be outside the global event horizon (which in the approximation we are using, is the horizon of the background black hole at $\rho^{D-3}=2m$).
@@ 5 Discussion
In our analysis we have shown that the black hole can be transformed into an ordinary quantum system containing stings and branes by letting it drop to a region close to a much bigger near extremal electrically charged black hole. We have not discussed how to extract information from this system, but since this is an ordinary quantum system, an observer falling with the rolling black hole should be able to study the quantum state of this system using standard methods and transmit the data to an asymptotic observer. The time scale of such experiments will be long due to the weakness of the string coupling and one might wonder whether the rolling black hole will fall into the background black hole before such experiments could be performed. Furthermore as the coupling becomes weaker as we approach the horizon of the background black hole the time scale of measurement becomes longer. The latter problem could be avoided by adding a small amount of magnetic charge to the background black hole so that the attractor value of the string coupling is given by $N^{-1/4}/\epsilon$ for some fixed small number $\epsilon$. In that case, even if the rolling black hole falls towards the horizon of the background black hole, the coupling does not decrease any further. Also one could in principle arrest the fall of the rolling black hole towards the background black hole using the classical gravitational field of other (accelerating) objects placed in appropriate regions of space-time (see footnote 1 for a discussion on this).. While we have not studied this problem in detail, we expect that this should be possible in principle as long as the system of which the rolling black hole is a part remains outside the horizon of the background black hole.
@@
Acknowledgement: This work was supported by the ICTS-Infosys Madhava Chair Professorship and the Department of Atomic Energy, Government of India, under project no. RTI4019.

### 09 Apr 2026 v2

@@ 1 Introduction
The main idea will be to use the correspondence principle formulated in [1, 2, 3], following earlier suggestions in [4, 5, 6, 7]. More recent perspective on this can be found in [8, 9, 10, 11]. This principle states that as we adiabatically change the string coupling from finite value to a sufficiently small value, a Schwarzschild black hole transforms to a highly excited elementary string state. This was generalized to a wide class of other charge carrying black holes, although at weak coupling these black holes correspond to systems of D-branes and fundamental strings instead of just fundamental strings.
@@
A similar mechanism for converting a black hole to an elementary string was considered in [11] with the help of large amplitude and large wavelength dilaton wave. However, as discussed in [14], there is an upper bound of order unity on the amplitude of a scalar wave in any dimensions, since for a wave of amplitude $a$ and wave-length $R$ measured in Planck units, the energy density is of order $(a/R)^{2}$ and the energy contained in a ball of size $R$ in $D$ space-time dimensions is of order $a^{2}R^{D-3}$. The associated Schwarzschild radius $a^{2/(D-3)}R$ will exceed $R$ for sufficiently large $a$, making the background collapse into a black hole. [^1] This is not a no go theorem since this argument does not take into account the full non-linear effects of gravity. However this shows that without having a more detailed analysis, we cannot trust the dilaton wave background to be horizon free. The background considered here does not suffer from this problem. Even though we use a large black hole to produce the desired background, the system that we study always remains outside the horizon of the large black hole and hence remains accessible to the asymptotic observer.
@@ 2.3 Varying dilaton background from charged black hole
Using these results we can also test the other conditions. For example using the lower bound on $C$ given in (2.32) we can verify that the mass of the background black hole, which is of order $C$, is much larger than the mass of the rolling black hole which is of order $N^{(D-3)/\{2(D-2)\}}$. Hence the backreaction of the rolling black hole on the background black hole geometry remains small. In particular, the causal structure of space-time remains unaffected and as long as the rolling black hole stays outside the horizon of the background black hole, given by the condition $\rho^{D-3}>2m$, it will be causally connected to the asymptotic observer. Also, it follows from (2.3) that for $\bar{\rho}\sim\bar{\rho}_{0}$, any scalar constructed from the fields and two derivatives scales as $\bar{\rho}_{0}^{-2/(D-2)}$. The only exception is $F_{\mu\nu}F^{\mu\nu}$, but the combination $e^{-4\phi/(D-2)}F_{\mu\nu}F^{\mu\nu}$ that appears in the Lagrangian density still scales as $\bar{\rho}_{0}^{-2/(D-2)}$. Hence for large $\bar{\rho}_{0}$, locally the rolling black hole experiences almost flat space-time. It also follows from this that a stationary observer in this background feels acceleration of order $\bar{\rho}_{0}^{-1/(D-2)}$. In the sense described in footnote 2, this also gives an order of magnitude estimate of the temperature experienced by the rolling black hole / string. From (2.31) we see that this is much smaller than the temperature $T_{bh}\sim N^{-1/\{2(D-2)\}}$ given in (2.14). Since (2.31) ensures that the effect of evaporation remains small during the transit, it now follows that the effect of absorption of the ambient radiation also remains small during the transit.
@@ 2.4 Fall from the asymptotic region
For $D\geq 5$ and large $N$, the upper bound is larger than the lower bound, and hence we can find $\bar{\rho}_{0}$ satisfying these conditions. For $D=4$, the $N$ dependence of the two sides are identical, but by taking $\eta$ to be small, we can still ensure that the upper bound is larger than the lower bound. Thus we see that it is possible for the rolling black hole to roll all the way from a region where the string coupling is small but $N$ independent to the region where it becomes an elementary string excitation. On the other hand, if the rolling black hole undergoes a controlled fall as discussed in footnote 2, maintaining its speed at order unity but not parametrically close to unity, then the lower bound on $\bar{\rho}_{0}$ will be given by the earlier bound $N^{1/2}$, and even for $D=4$ and finite $\eta$, the upper bound on $\bar{\rho}_{0}$ remains larger than the lower bound for large $N$.
Finally we can consider the fall from the asymptotic region $\rho\sim C^{1/(D-3)}$. In this case the transit time is of order $C^{1/(D-3)}$. Comparing this with the transit time $\bar{\rho}_{1}^{1/(D-2)}\sim C^{1/(D-3)}\eta^{2/\{(D-2)(D-3)\}}$ computed from (2.36) and (2.34), we see that this corresponds to setting $\eta\sim 1$ in the previous analysis. Also from (2.3) and (2.27) we see that the redshift factor $\gamma$ at a position $\rho\sim\bar{\rho}_{0}/C$ is now of order $(C\rho^{3-D})^{(D-3)/(D-2)}\sim N^{(D-3)/\{2(D-2)\}}$. Comparing this with (2.38) we again see that that this corresponds to setting $\eta\sim 1$ in the previous analysis. Thus for $D\geq 5$, even a free fall of the rolling black hole from the asymptotic region will convert a black hole to an elementary string state in a controlled fashion. For $D=4$ the adiabatic approximation will break down near the end of the roll, but this can be avoided by arresting the fall by external force as described in footnote 2 so that the rolling black hole does not become ultra-relativistic during the roll. In the extreme case, if we bring the rolling black hole at rest at $\bar{\rho}\sim\bar{\rho}_{0}\epsilon^{2/(D-3)}$ where it becomes an elementary string state, the proper acceleration felt by the rolling black hole will be of order $\bar{\rho}_{0}^{-1/(D-2)}\epsilon^{-2/\{(D-2)(D-3)\}}$. This remains small compared to the string scale as long as (2.30) holds. An observer in the rest frame of this black hole can study its quantum state and send the information back to the asymptotic observer since both the rolling black hole and the observer will be outside the global event horizon (which in the approximation we are using, is the horizon of the background black hole at $\rho^{D-3}=2m$).
@@ 5 Discussion
In our analysis we have shown that the black hole can be transformed into an ordinary quantum system containing stings and branes by letting it drop to a region close to a much bigger near extremal electrically charged black hole. We have not discussed how to extract information from this system, but since this is an ordinary quantum system, an observer falling with the rolling black hole should be able to study the quantum state of this system using standard methods and transmit the data to an asymptotic observer. The time scale of such experiments will be long due to the weakness of the string coupling and one might wonder whether the rolling black hole will fall into the background black hole before such experiments could be performed. Furthermore as the coupling becomes weaker as we approach the horizon of the background black hole the time scale of measurement becomes longer. The latter problem could be avoided by adding a small amount of magnetic charge to the background black hole so that the attractor value of the string coupling is given by $N^{-1/4}/\epsilon$ for some fixed small number $\epsilon$. In that case, even if the rolling black hole falls towards the horizon of the background black hole, the coupling does not decrease any further. Also one could in principle arrest the fall of the rolling black hole towards the background black hole using the classical gravitational field of other (accelerating) objects placed in appropriate regions of space-time (see footnote 2 for a discussion on this).. While we have not studied this problem in detail, we expect that this should be possible in principle as long as the system of which the rolling black hole is a part remains outside the horizon of the background black hole.
@@
Acknowledgement: I wish to thank Nejc Ceplak, Roberto Emparan, Andrea Puhm, Marija Tomasevic and Yoav Zigdon for useful communications. This work was supported by the ICTS-Infosys Madhava Chair Professorship and the Department of Atomic Energy, Government of India, under project no. RTI4019.